In one episode of the 1960s television sitcom, Gilligan’s Island, the famous “professor” constructed voltaic cells to use as substitutes for their radio’s dead batteries. A single D-cell has an emf of 1.5 V. Which scraps of metal from their damaged boat, the Minnow, could best be used to create a 1.5 V voltaic cell? Assume that coconuts make great beakers and that seawater is a terrific electrolyte.

   Iron/iron(II) (the anchor) -0.44 V

   Lead/lead(II) (fishing weights) -0.126 V

   Silver/silver(I) (Mrs. Howell’s brooch) -0.799 V

   Aluminum/aluminum(III) (the boat’s flagpole) -1.677 V
Answer #1

use the equation

E cell = E cathode - E anode

Lets try with lead cathode and silver anode

E cell = -0.126 - (-0.799) = 0.673

we are trying to get an Ecell = 1.5 V, lets try using the aluminum as an anode:

now use the iron as a cathode so:

E cell = -0.44 - (-1.677) = 1.237 V

Lets try lead cathode and aluminum anode:

E cell = -0.126 - ( - 1.677) = 1.551 V

using lead as a cathode and aluminum as an anode is the best choice.